OpenStudy (anonymous):

Find the perimeter of a parallelogram with one side that measures 38.0 cm and another that is 16 cm.

OpenStudy (michele_laino):

please the perimeter P of a parallelogram whose sides measure a and b is: P = 2*(a+b)

OpenStudy (idealist10):

P=2(s+b)

OpenStudy (idealist10):

P=2(38+16)=2(54)=108 Answer: 108 cm.

OpenStudy (michele_laino):

that's right!

OpenStudy (idealist10):

Remember, perimeter is the sum of all the sides of a polygon.